# RED-PSM: Regularization by Denoising of Partially Separable Models for Dynamic Imaging

### v3

@@ II Problem Statement
Our objective in dynamic tomography is to reconstruct the underlying dynamic object $\{f(\mathbf{x},t_{p})\}_{p=0}^{P-1}$ from the time-sequential projections in (1). The challenge is that because each projection belongs to a different object, the projections in (1) are inconsistent. Therefore, a conventional, e.g., filtered backprojection (FBP) reconstruction as for a static object results in significant reconstruction artifacts. This is to be expected, as the problem is severely ill-posed: an image with a $D$-pixels diameter requires more than $D$ projections for artifact-free reconstruction, whereas only one projection is available per time-frame in the time-sequential acquisition (1). Multiple dynamic tomography methods [32, 31, 33] group temporally neighboring projections and assume the object as static during their acquisition. However, this reduces the temporal resolution, and any violation of this assumption (as in the case of time-sequential sampling) leads to mismodeling in the data fidelity terms.
@@ III Partially Separable Models (PSM)
For spatio-temporal inverse problems such as dynamic MRI and tomography, the underlying object can be accurately represented using a partially-separable model (PSM), which effectively introduces a low-rank prior to the problem. For dynamic tomography, a PSM can represent the imaged object $f$, or its full set of projections $g$. In this paper, we use an object-domain PSM.
@@
This model facilitates interpretability by separating the spatial structure from the temporal dynamics. This expansion is dense in $\mathcal{L}_{2}$ [52], meaning that any finite energy object can be approximated arbitrarily well by such a model of sufficiently high order. Empirically, modest values of $K$ provide high accuracy in applications to MR cardiac imaging [12, 53, 54]. Theoretical analysis [10] shows that for a spatially bandlimited object undergoing a time-varying affine transformation (i.e, combination of time-varying translation, scaling, and rotation) of bounded magnitude, a low order PSM provides a good approximation.
As detailed later, we use a $d$-dimensional representation for the temporal functions, with $d\ll P$. Together with the PSM for the object this leads to a significant reduction in the number of representation parameters for a spatio-temporal object with a $D$-pixel diameter and $P$ temporal samples: from $\approx PD^{2}$ to $\approx KD^{2}+Kd$, with $K\ll P$. Because $d\ll D$, this corresponds to a factor of $P/K\gg 1$ compression of the representation, providing an effective parsimonious model for the spatio-temporal object $f$.
@@ IV-A Variational Formulation
The first term is the data fidelity term measuring the fit between available undersampled measurements $g_{t}$ of the true object and the measurements obtained from the estimated object $f=\Lambda\Psi^{T}\in\mathbb{R}^{N^{2}\times P}$. The second term with weight $\lambda>0$ is a spatial regularizer injecting relevant spatial prior to the problem. It is applied to the PSM $\Lambda\Psi^{T}$ column by column, that is, to individual temporal image frames. The last two terms with weight $\xi>0$ prevent the bilinear factors from growing without limit.
@@ IV-B Incorporating Regularization by Denoising
where $D_{\phi}$ is the denoiser. This results in Algorithm 2, which requires only single use of the denoiser per iteration of ADMM. Furthermore, as in Algorithm 1, the modified Line 4 in Algorithm 2 can be performed in parallel for all $t$.
@@ IV-D Computational Cost
Space (memory requirements) and time (operation count) complexities for two variants of the RED-PSM bilinear ADMM Algorithm are shown in Table I. The operation counts are for a single outer iteration. As the number of outer iterations typically has a weak dependence on the size of the problem, the scaling shown tends to determine the run time. Furthermore, thanks to its structure, the algorithm also offers many opportunities for easy parallelization, so that actual runtime can be proportionally reduced by allocating greater computational resources. See the Supplementary Material Section A-B for the detailed analysis of computational requirements.
@@ V Convergence Analysis
In this section, we follow an approach similar to recent work on ADMM for a bilinear model [62] to analyze convergence. We show that under mild technical conditions, the objective in Algorithm 1 is guaranteed to converge (with increasing number $I$ of iterations) to a value corresponding to a stationary point of the Lagrangian, that is, satisfying the necessary conditions for the first order optimality.
In practice, Algorithm 2 with the efficient $f$ step version, which we implemented and used in the experiments reported in Section VI, has better run times, and rapid empirical convergence. However, its analysis requires additional steps, which are not particularly illuminating. Therefore, following the tradition in related analyses [58, 63], we focus on the analysis of the nominal Algorithm 1.
The Frobenius norm penalties on basis functions result, as shown in Appendix A, in the boundedness of the iterates $(f^{(i)},\Lambda^{(i)},\Psi^{(i)})$.
Different to [62], our objective function has factorized terms in bilinear form. Also, our analysis includes a proof (similar to [64]) of the boundedness of the iterates to justify the existence of points of accumulation of the iterate sequence, which appears to have been inadvertently left out in [62].
@@
Moreover, we assume the gradient rule of RED (8). This is shown to hold true when the denoiser $D$ is assumed locally homogeneous (LH) [17] and Jacobian symmetric (JS) [24]. In [17], it was shown that several popular denoisers are practically LH, and [24] showed that for (8) to hold for the explicit regularizer $\rho$ in (4), JS is necessary.
In our analysis, we only require the gradient rule to hold, rather than the explicit form (4) of the regularizer. Even for denoisers without LH and JS, the gradient rule can be used to define the fixed point iteration in the algorithm [24], and, more recent works on using denoisers for regularization [65, 66, 67, 68, 69] (in particular [67]) shows how to learn a denoiser jointly with the gradient so that the gradient rule is satisfied.
@@
Suppose that the denoiser $D_{\phi}$ satisfies the gradient rule in (8), is Lipschitz continuous with some Lipschitz constant $L_{D}$, and strongly passive.
@@
The proof of Theorem 1 is provided in Appendix A.
Regarding convergence to the globally optimal solution, more can be said. Problem (IV-A) is non-convex even if $\bar{\rho}$ is convex, because of the product between unknowns $\Lambda$ and $\Psi$. However, similar to [63], thanks to the inclusion of the Frobenius norms of the factors $\Lambda$ and $\Psi$ in the bilinear form, when $\bar{\rho}$ is convex (i.e., when LH and JS assumptions hold), the global minimum $\hat{f}=\hat{\Lambda}\hat{\Psi}^{T}$ of (IV-A) can be shown to coincide with the global minimum in
@@
Problem (16) is non-convex too, because of the rank constraint. However, consider the convex version of this problem, without the rank constraint, and denote its solution by $\hat{f}^{\mathrm{Convex}}$. If $\mathrm{rank}(\hat{f}^{\mathrm{Convex}})\leq K$ then the low-rank constraint in Problem (16) is not active, and $\hat{f}^{\mathrm{Convex}}=\hat{f}$. It then follows [63] that the optimality conditions of the convex problem can be used to assess the global optimality of the solution $\hat{f}~{}=~{}\hat{\Lambda}\hat{\Psi}^{T}$ produced by the algorithm.  [^3] The work in [63] credits an equivalence between minimizing the sum of the Frobenius norm of the factors and the minimization of the nuclear norm of the product to [70]. However, the analysis in [70] is not directly applicable to our scenario. It would require the data fidelity $R_{\theta_{t}}\Lambda\Psi^{T}e_{t}=g_{t}$ as a hard constraint, includes only the Frobenius norm penalties in the minimization objective without additional explicit regularization, and requires the measurement operator to satisfy a Restricted Isometry Property (RIP). We are not aware of an analysis showing RIP for the measurement models in this paper. These aspects prevent inferring an equivalence between our problem and nuclear norm minimization directly from [70].
@@ VI-A Datasets
In a conference submission [75] citing this work, the method has been applied (without describing the algorithm or its theory) to additional tomographic scenarios of materials compressed under load, further confirming the advantage of the proposed method and algorithm over the compared methods.
@@
VI-B Baseline Methods
[…]
TD-DIP [18]: TD-DIP is a recent method based on the Deep Image Prior (DIP) approach. [^4] It would be interesting to include yet another baseline (also developed for dMRI) – the DIP-based PSM approach [45]. However, as an implementation of this method was unavailable, and due to potential issues (specific initialization and framework that use other MRI algorithms) with replicating its performance and adapting to our CT problem, we were unable to do so. It uses a mapping network $M_{\alpha}$ and a generative model $N_{\beta}$ in cascade to obtain the estimated object at each time instant $f_{t}$ from fixed and handcrafted latent representations $\chi_{t}$. Because TD-DIP was originally proposed for dynamic MRI, we modified the objective minimally for dynamic tomography as
@@ VI-B Baseline Methods
For the comparisons in this work, identical mapping network and generator architectures, latent representation dimensionality, optimizer, learning rate, and decay schemes are used as in the available implementation [76]. The original work focuses on the beating heart problem and thus proposes a helix-shaped manifold for $\chi_{t}$ with number of cycles equal to the number of heartbeats during measurement acquisition. Since we do not have a repetition assumption for the motions of the walnut and compressed object, we use a linear manifold as explained in the original paper [18]. Thus, for clarity, in Section VI-D the method is sometimes denoted as “TD-DIP (L)”.
@@ VI-C Experimental Settings
Denoiser and training. Each convolutional layer in the denoiser network is followed by a ReLU nonlinearity except for the final single-channel output layer. We tested both direct and residual DnCNN denoisers, where the former predicts the denoised image and the latter estimates the noise from the input. We use the denoiser type that performs better for each object, but the differences are minor. Further architectural details for the denoisers in our experiments are provided in Table VII in the Supplementary Material. We use three pre-trained denoisers, one for each of the three object types. In each case, the same pre-trained denoiser was used for all values of $P$.
The upper limit for noise level used in training the denoiser was set to $\sigma_{\max}=5\cdot 10^{-2}$. For the dynamic walnut object, the denoiser $D_{\phi}$ is trained on the central 200 axial, 200 sagittal, and 200 coronal slices of another static walnut CT reconstruction downsampled to size 128$\times$128. For the compressed object, axial slices of pre-compression and post-compression static versions of the object, containing 462 slices in total, are used to train $D_{\phi}$. For the cardiac MRI setting, the denoiser was trained on the static MRI training slices of the ACDC dataset [78]. For all datasets, $D_{\phi}$ is trained for 500 epochs using the Adam optimizer with a learning rate of $5\cdot 10^{-3}$. We test both direct and residual DnCNN denoisers where the former predicts the denoised image and the latter estimates the noise from the input.
Architectural details for denoisers in our experiments are provided in Table VII in Supplementary Material A-A. We use a fixed pre-trained denoiser for all $P$ for the same object type.
[…]
Initialization. Unless stated otherwise, the spatial and temporal basis functions are initialized using the SVD truncated to the rank of the dynamic object estimate produced by a recent projection-domain PSM-based method “ProSep” [10]. If the ProSep estimate has rank smaller than $K$, the remaining basis functions are initialized as 0.
Otherwise, all spatial basis functions are initialized as 0 and the temporal latent representations $z_{k}$ are initialized randomly as $z_{k}\sim\mathcal{N}(0,I)$.
@@ VI-D1 Reconstruction accuracies for different $P$
RED-PSM is compared in Figure 4 and Table II with PSM-TV-S, PSM-TV-ST and TD-DIP (using, for both a-periodic objects, a linear latent representation). Parameter configurations for the experiments are provided in Table V in Supplementary Material A-A. While Figure 4 facilitates the assessment and comparison of trends in the metrics for different $P$ for the various methods, Table II provides the detail for a more precise quantitative comparison. In fact, since the plotted range of metrics such as PSNR in Figure 4 is very large due to varying $P$, Table II emphasizes important differences between methods for the same $P$.
@@
We would like to emphasize the significance of these reconstruction quality improvements by comparing them to some representative examples. TD-DIP reports up to 3 dB PSNR and 0.005 SSIM improvement with respect to an older dMRI method in a single scenario. While providing comparable PSNR improvements (2.4 dB), we are able to provide four times larger SSIM (0.02) improvements relative to TD-DIP itself. Other recent dMRI [44] and dynamic passive acoustic tomography [46] methods improve 1.5 dB, and at most 1 dB and often no improvement over their respective benchmarks, again in single imaging scenario for each method.
@@
Figure 5 compares the reconstructions for both objects at two different values of $t$ for $P=256$. As expected, PSM-TV-S performs the worst among the compared methods and provides, for both objects, blurry reconstructions lacking finer details. The TD-DIP reconstructions improve somewhat over PSM-TV-S and PSM-TV-ST, but contain visible noise-like artifacts on the piecewise constant regions of the walnut object, which are alleviated by RED-PSM. This is manifested also in the absolute difference figures, with error for TD-DIP distributed throughout the interior regions of the walnut. Also, around the shell of the walnut, RED-PSM is further able to preserve sharper details. For the compressed material, in comparison to TD-DIP, RED-PSM shows reduced absolute error almost uniformly over the object. This difference is more prominent around the highly dynamic regions of the object, emphasizing the advantage of the proposed method. For a better understanding of RED-PSM results, we also display the reconstructed spatial and temporal basis $\Lambda$ and $\Psi$ for a time-varying walnut setting in Supplementary Material A-C.

### v4

@@ II Problem Statement
Our objective in dynamic tomography is to reconstruct the underlying dynamic object $\{f(\mathbf{x},t_{p})\}_{p=0}^{P-1}$ from the time-sequential projections in (1). The challenge is that because each projection belongs to a different object, the projections in (1) are inconsistent. Therefore, a conventional, e.g., filtered backprojection (FBP) reconstruction as for a static object results in significant reconstruction artifacts. This is to be expected, as the problem is severely ill-posed: an image with a $D$-pixels diameter requires more than $D$ projections for artifact-free reconstruction, whereas only one projection is available per time-frame in the time-sequential acquisition (1). Several dynamic tomography methods [89, 79, 100] group temporally neighboring projections and assume the object is static during their acquisition. However, this reduces the temporal resolution, and any violation of this assumption (as in the case of time-sequential sampling) leads to mismodeling in the data fidelity terms.
@@ III Partially Separable Models (PSM)
For spatio-temporal inverse problems such as dynamic MRI and tomography, the underlying object can be accurately represented using a partially-separable model (PSM), which effectively introduces a factorized low-rank prior to the problem. For dynamic tomography, a PSM can represent the imaged object $f$, or its full set of projections $g$. In this paper we use an object-domain PSM.
@@
This model facilitates interpretability by separating the spatial structure from the temporal dynamics. Empirically, modest values of $K$ provide high accuracy in applications to MR cardiac imaging [10, 107, 108]. Theoretical analysis [94] shows that for a spatially bandlimited object undergoing a time-varying affine transformation (i.e, combination of time-varying translation, scaling, and rotation) of bounded magnitude, a low order PSM provides a good approximation.
As detailed later, as a standard choice in modeling temporal functions, we use a $d$-dimensional representation with $d\ll P$. A similar idea was used in [11] to represent the low temporal bandwidth of cardiac motion in dMRI. Together with the PSM for the object this leads to a significant reduction in the number of representation parameters for a spatio-temporal object with a $D$-pixel diameter and $P$ temporal samples: from $\approx PD^{2}$ to $\approx KD^{2}+Kd$, with $K\ll P$. Because $d\ll D$, this corresponds to a factor of $P/K\gg 1$ compression of the representation, providing an effective parsimonious model for the spatio-temporal object $f$.
@@ IV-A Variational Formulation
The first term is the data fidelity term measuring the fit between available undersampled measurements $g_{t}$ of the true object and the measurements obtained from the estimated object $f=\Lambda\Psi^{T}\in\mathbb{R}^{N^{2}\times P}$. The second term with weight $\lambda>0$ is a spatial regularizer injecting relevant spatial prior to the problem. It is applied to the PSM $\Lambda\Psi^{T}$ column by column, that is, to individual temporal image frames. The last two terms with weight $\xi>0$ prevent the bilinear factors from growing without limit.  [^2] Inclusion of these Frobenius norm terms also happens to lead to connections with nuclear norm, but, as discussed in detail in Section V, our problem formulation (even if the RED regularizer is disregarded) is not equivalent to a formulation encouraging a low-rank solution by a nuclear norm or other Schatten norm penalty such as used in [44, 46, 47]. Instead, the factorized form $f=\Lambda\Psi^{T}$ (as in other PSM-based matrix recovery works such as [10, 11]) introduces this rank constraint explicitly.
@@ IV-B Incorporating Regularization by Denoising
where $D_{\phi}$ is the denoiser. This results in Algorithm 2, which requires only a single use of the denoiser per iteration of ADMM. Furthermore, as in Algorithm 1, the modified Line 4 in Algorithm 2 can be performed in parallel for all $t$.
@@ IV-D Computational Cost
Space (memory requirements) and time (operation count) complexities for two variants of the RED-PSM bilinear ADMM algorithm are shown in Table I. The operation counts are for a single outer iteration. As the number of outer iterations typically has a weak dependence on the size of the problem, the scaling shown tends to determine the run time. Furthermore, thanks to its structure, the algorithm also offers many opportunities for easy parallelization, so that actual runtime can be proportionally reduced by allocating greater computational resources. See the Supplementary Material Section VII-B for the detailed analysis of computational requirements.
@@ V Convergence Analysis
In this section, we follow an approach similar to recent work on ADMM for a bilinear model [116] to analyze convergence. We show that under mild technical conditions, the objective in Algorithm 1 is guaranteed to converge (with increasing number $I$ of iterations) to a value corresponding to a stationary point of the Lagrangian, that is, satisfying the necessary conditions for first order optimality.
In practice, Algorithm 2 with the efficient $f$ step version, which we implemented and used in the experiments reported in Section VI, has better run times, and rapid empirical convergence. However, its analysis requires additional steps, which are not particularly illuminating. Therefore we focus on the analysis of the nominal Algorithm 1.
In spite of the similarity, at a high level, to the problem and analysis in [116], our problem formulation and algorithm differ in several aspects, which require modifications in the analysis and proof. In particular, different to [116], where the bilinear form only appears in the constraint, the bilinear form appears both in our objective function and constraint. To satisfy strong convexity for the respective subproblems, our objective also uses the Frobenius norm terms for the PSM factors, instead of the interim proximal terms in [116]. Moreover, to allow the efficient $f$-step in Algorithm 2, we use a different order of updates than [116]. Importantly, we account for the explicit RED regularization and its required properties in the proof of convergence. Our analysis includes a proof (similar to [117]) of the boundedness of the iterates to justify the existence of points of accumulation of the iterate sequence, which appears to have been inadvertently left out in [116]. However, different to [117], which has objectives with linear constraints, our objective (IV-A) has a bilinear constraint.
@@
Moreover, we assume the gradient rule of RED (8). This is shown to hold true when the denoiser $D$ is assumed locally homogeneous (LH) [13] and Jacobian symmetric (JS) [20]. It was shown [13] that several popular denoisers are practically LH, and that for (8) to hold for the explicit regularizer $\rho$ in (4), JS is necessary [20].
In our analysis, we only require the gradient rule (8) to hold, rather than the explicit form (4) of the regularizer. Even for denoisers without LH and JS, the gradient rule can be used to define the fixed point iteration in the algorithm [20], and, more recent works on using denoisers for regularization [118, 119, 120, 121, 122] (in particular [120]) shows how to learn a denoiser jointly with the gradient so that the gradient rule is satisfied.
@@
Suppose that the denoiser $D_{\phi}$ satisfies the gradient rule in (8), and is Lipschitz continuous with some Lipschitz constant $L_{D}$, and strongly passive.
@@
The proof of Theorem 1 is provided in the supplementary material Section D.
Regarding convergence to the globally optimal solution, more can be said. Problem (IV-A) is non-convex even if $\bar{\rho}$ is convex, because of the product between unknowns $\Lambda$ and $\Psi$. However, similar to Theorem 7.1 of [123] (which generalizes arguments in [124],[125]), thanks to the inclusion of the Frobenius norms of the factors $\Lambda$ and $\Psi$ in the bilinear form, when $\bar{\rho}$ is convex (i.e., when the gradient rule (8) is assumed to hold), the global minimum $\hat{f}=\hat{\Lambda}\hat{\Psi}^{T}$ of (IV-A) can be shown to coincide with the global minimum in
@@
Importantly, Problem (16) is non-convex too, because of the rank constraint. In other words, it is not equivalent to a conventional nuclear norm penalized problem. Hence, although the Frobenius norm penalties on the PSM factors in our problem formulation (IV-A) lead to a connection to nuclear norm, Problem (IV-A) is not equivalent to a formulation encouraging a low-rank solution by a nuclear norm penalty.
However, returning to the question of determining global optimality of a candidate solution to (16), consider the convex version of Problem (16), without the rank constraint, and denote its solution by $\hat{f}^{\mathrm{Convex}}$. If it so happens that $\mathrm{rank}(\hat{f}^{\mathrm{Convex}})\leq K$, then the low-rank constraint in Problem (16) is not active, and $\hat{f}^{\mathrm{Convex}}=\hat{f}$. It then follows [123] that the optimality conditions of the convex problem (without the rank constraint) can be used to assess the global optimality of a candidate solution $\hat{f}~{}=~{}\hat{\Lambda}\hat{\Psi}^{T}$ produced by the algorithm solving (IV-A).
@@ VI-A Datasets
In a conference submission [130] citing this work, the method has been applied to additional tomographic scenarios of materials compressed under load, further confirming the advantage of the proposed method and algorithm over the compared methods.
@@
VI-B Comparison Benchmarks
PSM-TV: Similar to the proposed approach, this algorithm also uses a PSM to represent the object, but instead of the RED regularizer, the regularization penalizes the discrete 2D total variation of the temporal frames of $f$ at each time instant. To this end, the constraint $f=\Lambda\Psi^{T}$ is implemented by substitution into the objective in (5), and the definition of $\rho$ is changed to $\rho(\cdot)=\text{TV}(\cdot)$, and the rest of the objective is kept the same. We consider spatial (PSM-TV-S) and spatiotemporal (PSM-TV-ST) alternatives of TV. The former computes TV only spatially in a single frame at each $t$ whereas the latter also computes differences between temporally adjacent pixels at $t-1$ and $t+1$. The unconstrained problem is then solved for $\{\hat{\Lambda},\hat{\Psi}\}$ (using Adam optimizer in Pytorch). Finally, the estimated object is obtained as $\hat{f}=\hat{\Lambda}\hat{\Psi}^{T}$.
TD-DIP [14]: TD-DIP is a recent method based on the Deep Image Prior (DIP) approach.  [^4] It would be interesting to include yet another comparison benchmark (also developed for dMRI) – the DIP-based PSM approach [74]. However, as an implementation of this method was unavailable, and due to potential issues with replicating its performance and adapting to our CT problem (specific initialization and framework that use other MRI algorithms), we were unable to do so. It uses a mapping network $M_{\alpha}$ and a generative model $N_{\beta}$ in cascade to obtain the estimated object at each time instant $f_{t}$ from fixed and handcrafted latent representations $\chi_{t}$. Because TD-DIP was originally proposed for dynamic MRI, we modified the objective minimally for dynamic tomography as
@@ VI-B Comparison Benchmarks
For the comparisons in this work, the mapping network and generator architectures, latent representation dimensionality, optimizer, learning rate, and decay schemes are are identical to those in the available implementation [131]. The original work focuses on the beating heart problem and thus proposes a helix-shaped manifold for $\chi_{t}$ with number of cycles equal to the number of heartbeats during measurement acquisition. Since we do not have a repetition assumption for the motions of the walnut and compressed object, we use a linear manifold as explained in the original paper [14]. Thus, for clarity, in Section VI-D the method is sometimes denoted as “TD-DIP (L)”.
@@ VI-C Experimental Settings
Denoiser and training. Each convolutional layer in the denoiser network is followed by a ReLU nonlinearity except for the final single-channel output layer. We tested both direct and residual DnCNN denoisers, where the former predicts the denoised image and the latter estimates the noise from the input. We use the denoiser type that performs better for each object, but the differences are minor. Further architectural details for the denoisers in our experiments are in Table VII in the Supplementary Material. We use three pre-trained denoisers, one for each of the three object types. In each case, the same pre-trained denoiser was used for all values of $P$.
The upper limit for noise level used in training the denoiser was set to $\sigma_{\max}=5\cdot 10^{-2}$. For the dynamic walnut object, the denoiser $D_{\phi}$ is trained on the central 200 axial, 200 sagittal, and 200 coronal slices of another static walnut CT reconstruction downsampled to size 128$\times$128. For the compressed object, axial slices of pre-compression and post-compression static versions of the object, containing 462 slices in total, are used to train $D_{\phi}$. For the cardiac MRI setting, the denoiser was trained on the static MRI training slices of the ACDC dataset [133]. For all datasets, $D_{\phi}$ is trained for 500 epochs using the Adam optimizer with a learning rate of $5\cdot 10^{-3}$. As mentioned above, we evaluate both direct and residual DnCNN denoisers.
Temporal Basis. In compressed material and cardiac dMRI data experiments, we use the parametrization $\Psi=UZ$ with a fixed basis $U$ that corresponds to a cubic spline interpolator, and for the warped walnut we use DCT-II, to interpolate the low-dimensional temporal representation $Z$ to $\Psi$.
Initialization. Unless stated otherwise, the spatial and temporal basis functions are initialized using the SVD truncated to the rank of the dynamic object estimate produced by a recent projection-domain PSM-based method “ProSep” [94]. If the ProSep estimate has rank smaller than $K$, the remaining basis functions are initialized as 0. Otherwise, all spatial basis functions are initialized as 0 and the temporal latent representations $z_{k}$ are initialized randomly as $z_{k}\sim\mathcal{N}(0,I)$.
@@ VI-D1 Reconstruction accuracies for different $P$
RED-PSM is compared in Figure 4 and Table II with PSM-TV-S, PSM-TV-ST and TD-DIP (using, for both a-periodic objects, a linear latent representation). Parameter configurations for the experiments are provided in Table V in Supplementary Material A. While Figure 4 facilitates the assessment and comparison of trends in the metrics for different $P$ for the various methods, Table II provides the detail for a more precise quantitative comparison. In fact, since the plotted range of metrics such as PSNR in Figure 4 is very large due to varying $P$, Table II emphasizes important differences between methods for the same $P$.
@@
We would like to emphasize the significance of these reconstruction quality improvements by comparing them to some representative examples. TD-DIP reports up to 3 dB PSNR and 0.005 SSIM improvement with respect to an older dMRI method in a single scenario. While providing comparable PSNR improvements (2.4 dB), we are able to provide four times larger SSIM (0.02) improvements relative to TD-DIP itself. Other recent dMRI [73] and dynamic photo-acoustic tomography [102] methods improve 1.5 dB, and at most 1 dB and often no improvement over their respective benchmarks, again in single imaging scenario for each method.
@@
Figure 5 compares the reconstructions for both objects at two different values of $t$ for $P=256$. As expected, PSM-TV-S performs the worst among the compared methods and provides, for both objects, blurry reconstructions lacking finer details. The TD-DIP reconstructions improve somewhat over PSM-TV-S and PSM-TV-ST, but contain visible noise-like artifacts on the piecewise constant regions of the walnut object, which are alleviated by RED-PSM. This is manifested also in the absolute difference figures, with error for TD-DIP distributed throughout the interior regions of the walnut. Also, around the shell of the walnut, RED-PSM is further able to preserve sharper details. For the compressed material, in comparison to TD-DIP, RED-PSM shows reduced absolute error almost uniformly over the object. This difference is more prominent around the highly dynamic regions of the object, emphasizing the advantage of the proposed method. For a better understanding of RED-PSM results, we also display the reconstructed spatial and temporal basis $\Lambda$ and $\Psi$ for the time-varying walnut scenario in Supplementary Material C.